undeadknight26 (undeadknight26):

Just to be sure...How would you solve this? 1/2 + 1/3 =1/x

OpenStudy (anonymous):

make the denominators the same for the left side, least common denominator is what

OpenStudy (anonymous):

5/6=1/x so inverse we have: x=6/5

OpenStudy (shaadykhamlichi):

1/2+1/3=1/x 3/6+2/6=1/x 5/6=1/x 5x=6 x=6/5

OpenStudy (shaadykhamlichi):

is it helpful

undeadknight26 (undeadknight26):

Yes thank you guys :D

OpenStudy (anonymous):

Hello Tariq Rahaman.

OpenStudy (anonymous):

I will be contacting your parents for a withdrawal from OCVS.

OpenStudy (anonymous):

lol did you get scared?????

undeadknight26 (undeadknight26):

@Mr.Booker_OCVS I will kick your ash ._.

OpenStudy (shaadykhamlichi):

can i have a medal and a fan

undeadknight26 (undeadknight26):

Sure why not :)

OpenStudy (shaadykhamlichi):

ok

OpenStudy (anonymous):

lol